painting on a piece of canvas that has an area 180 square inches. The length of the painting is 1 1\4 times the width. What are the dimensions of the painting.
Answer by isma95fb(6)

let length of painting = L and width = W
from area = length x width
180 inch^2 = L x W
but L = 5/4 x W
180 inch^2 = 5/4 x W x W
W^2 = (4 x 180)/5
W = 12
but L = 5/4 x W
L = 5/4 x 12
L = 15
hence the dimensions are 15 inches long and 12 inches wide